# Career Modeling with Missing Data and Traces

### v2

@@ Abstract
Many social scientists study the career trajectories of populations of interest, such as economic and administrative elites. However, data to document such processes are rarely completely available, which motivates the adoption of inference tools that can account for large numbers of missing values. Taking the example of public-private paths of elite civil servants in France, we introduce binary Markov switching models to perform Bayesian data augmentation. Our procedure relies on two data sources: (1) detailed observations of a small number of individual trajectories, and (2) less informative “traces” left by all individuals, which we model for imputation of missing data. An advantage of this model class is that it maintains the properties of hidden Markov models and enables a tailored sampler to target the posterior, while allowing for varying parameters across individuals and time. We provide two applied studies which demonstrate this can be used to properly test substantive hypotheses, and expand the social scientific literature in various ways. We notably show that the rate at which ENA graduates exit the French public sector has not increased since 1990, but that the rate at which they come back has increased.
Keywords: Bayesian Markov Switching Model, Career Modeling, Public-Private Paths, Civil Servants, Missing Values
@@ 1 Introduction
Many important questions in the social sciences require understanding career trajectories of certain populations. This is an important aspect when describing political (Jäckle and Kerby, 2018), economic (Dudouet et al., 2010), academic (Angermuller and Blanchard, 2023) and administrative (Rouban, 2014) elites, but also when studying specific professions like musicians (Abbott and Hrycak, 1990), journalists (Machut and Bastin, 2023), or even athletes (Li et al., 2018). Studying careers is seen as interesting in its own right, but also to describe broader social mechanisms such as lobbying and revolving doors (Blanes i Vidal et al., 2012) or the spatial structuring of finance (Bühlmann et al., 2024).
@@
However, a key problem in career modeling is data missingness. It is very rare that all data relevant to the career of a certain population are available in a public record or can even be collected because of resource constraints, which is problematic for two reasons. First, for a given defined population, inference can be made more difficult and less powerful if a majority of individuals cannot be observed. This reduces the breadth of questions that can be answered using career modeling, as it makes the testing of interesting interaction effects impossible in practice. Second, researchers often choose to only study populations that can be entirely observed. This forces them to study the groups that are the most visible, such as top political elites; it thus constrains the breadth of questions that can be addressed using career modeling.
Our article has two aims. First, we propose a general methodology employing hidden Markov models with covariates and autoregressive components, and showcase, using plausible synthetic data examples, why they are required for such a modeling task while remaining simple enough for substantive hypothesis testing. The procedure relies on (i) a fraction of the population being well-observed using a first (possibly hand collected) data source, and (ii) the usage of partial “traces” from a digital or administrative record that are modeled to input the missing information for the remaining majority of individuals. Second, we provide two applied studies showcasing the usefulness of these models, which expand the social scientific literature of administrative elites in meaningful ways.
@@
Our leading example will be the study of public-private paths of elite civil servants in France. Paths of civil servants into the private sector are seen as important both by the general public and in the social science literature, where many publications that quantitatively study French high civil servants tackle the issue directly (Bouzidi et al., 2010; Kolopp, 2021; Rouban, 2014). Public-private paths broadly understood for policy-relevant positions are sometimes described using the term “revolving doors”, because these back-and-forth switches resemble a physical revolving door (Seabrooke and Tsingou, 2021; France and Vauchez, 2017), but other scholars define revolving doors to be a much more precise and limited in scope phenomenon. Unassuming about this definitional debate, we denote this problem as the study of public-private paths of French elite civil servants, and acknowledge that our article falls in the first category of works.
The phenomenon of public-private paths of civil servants demonstrates well the usefulness of the general method because it shares many difficulties with other settings in career analysis. First, it is relatively easy to describe the phenomenon for a portion of elite civil servants, by collecting data from digital networking websites such as LinkedIn on which they are registered. However, a majority of them do not have a networking digital profile, and further, it is not always the case that a profile is fully completed from the starting point of one’s career. Second, most civil servants, to various degrees, leave “traces” of their activity, when they sign legal documents, or obtain important promotions or bonuses. We showcase that it is possible to reconstruct trajectories of large civil servant populations and draw interpretations that contribute to the social scientific literature. In addition, our population is voluntarily more diverse and an order of magnitude larger compared to usual prosopography studies of the domain, to demonstrate that our methodology is well suited for analyses that are also targeted at less visible, larger professional groups.
We introduce and discuss in Section 2 the general statistical framework for career modeling with partially observed data, then apply it on real data to conduct two studies on French elite civil servants in Section 3. Section 4 concludes both on the general methodology and on the substantive learning exercise on civil servants.
2 General methodology
2.1 Problem setting
2.1.1 Detailed and trace data sources
@@
The core of our approach stems from the tension between data quality and quantity. In the majority of cases, it is difficult to obtain data that exhibit both high precision to document career trajectories, and high coverage. However, there often exists at least two data sources with opposed strengths and weaknesses:
  • Detailed data source. One data source exhibits low coverage but is very informative on the individuals it documents. This usually corresponds to data collected by hand for a small number of individuals, but can also be taken automatically from digital or administrative sources, such as professional networking websites.
@@
  • Trace data source. Another data source exhibits high coverage but is not very informative taken on its own and requires statistical elaboration for analysis. This concerns larger administrative or digital records, which we will describe as collections of “traces”: we know that individuals in a certain career position should regularly leave traces of their activity in these records, but the rate at which this happens varies across people and time.
Let us provide several examples taken from the introduction. When studying journalists’ affiliations, we know that some of them have digital profiles on networking websites (detailed data source), or we can spend some time hand-crafting a career database on a subpopulation; but we also know that journalists regularly publish newspaper articles with their name and affiliation (trace data source). We can apply the same logic to academic researchers, even if the rate at which they publish is lower: modeling the process by which individuals emit traces needs to be informed by data, and this motivates our mathematical framework. For board members in the economic elite, one might hand-construct a dataset (detailed data source); but we might also find their names with affiliation in documents at the end of every board meeting (trace data source), which might vary in frequency from company to period. Our leading example will be the case of elite civil servants in the public and private sectors: a good portion of them have a digital presence (detailed data source); but for many groups, we know they leave traces of their activity in administrative records when they get certain promotions or take certain decisions in the public sector (trace data source). For all the examples cited in the introduction, one could find credible data sources that would match this setting.
2.1.2 Formalization
@@
We now turn to the formalization of this setting. We aim to study the career trajectory of $N$ individuals $i\in\{1,...,N\}$ in discrete time $t\in\{1,...,T\}$. We denote the career position of individual $i$ in time $t$ as $X_{i,t}\in\mathcal{X}$, with $\mathcal{X}=\{x_{1},...,x_{K}\}$ a known finite state space of size $K$. In our application, $\mathcal{X}$ is equal to $\{0,1\}$, which corresponds to public and private sectors, but this can describe many other settings: position in a given newsroom or media type for a journalist, stage of the career for an academic, company for a corporate manager… As we will use a model in the family of Hidden Markov Models (HMMs), we expect that choosing a higher cardinality for $\mathcal{X}$ will incur further assumptions so that the parameters governing the behavior remain of moderate effective dimension. The ensemble of all trajectories at all time points is denoted $\textbf{X}=(X_{i,t})_{i,t}$. Observe that time here is discrete: our techniques may be generalized with continuous time, but we do not find a sufficient interest given the prohibitive cost that would incur. In our applications, $t$ corresponds to a six-month period; other choices could be reasonable depending on the substantive research question at hand. We assume that X is partially observed, thanks to the detailed data source: for some (but not all) values of $i$ and $t$, we observe the value of $X_{i,t}$.
@@
We now need to connect it to information found in the trace data source. The trace data source is described using a second set of time series, $\textbf{Y}=(Y_{i,t})_{i,t}$, which take value $Y_{i,t}\in\mathcal{Y}$ for each $i$ at time $t$. The state space $\mathcal{Y}$ may here be either discrete or continuous, one-dimensional or multi-dimensional. In our application, it will simply be a binary variable describing whether an administrative trace can be found for individual $i$ at time $t$. This could correspond to the number of articles published by a journalist or an academic, an encoding of an athlete’s participation in different competitions, or a binary encoding on whether there was a document including the name of a board member in the given period. Unlike X, we assume that Y is fully observed.
@@
Lastly, this setting is described with covariates $\textbf{W}=(W_{i,t})_{i,t}$ which are of interest to the researchers or which are deemed important to control for, and which may vary across individuals and time. A typical example will be to include the sex of individuals, their age, or their initial diplomas, to see if career behavior differs across these characterics. Importantly, these covariates will be used both to model the behavior of career paths X, and the relationship between paths X and traces Y, and so may include both substantive and nuisance characteristics. We will show it is possible and important to include transformations of lagged values of $Y_{i,t}$ into $W_{i,t}$ when describing the relationship between X and Y.
2.2 Model building with Markov switching models and inference
2.2.1 Related works and inspirations
Our modeling approach is inspired by duration analysis models and their connection to sequence analysis, and by recent developments using Markov switching models in realistic applied settings.
When working with social sequence data, many questions are fundamentally tied to duration. For example, duration analysis models have been applied the durability of political elites such as ministers (King et al., 1990; Bienen and Van de Walle, 1991). In addition, many applications find interest in the duration individuals may stay in more than one type of position: Machut and Bastin (2023) are interested in the career instability of journalists and how transitions across steps of the careers are accelerating with time; Angermuller and Blanchard (2023) are interested in the mobility of academic researchers across statuses, institutions, and countries; and we are interested in both the time elite civil servants spend in the public sector and the time they take before coming back. In our specific application, we complement the scholarship on the tenure of administrative elites using traditional duration analysis (Fleischer, 2016) since we also consider the time taken before coming back, and similar generalizations for the other examples can be provided by modeling as a whole trajectories X instead of only focusing on a single duration problem.
@@ 2.2.1 Related works and inspirations
One traditional solution is to study discrete trajectories as Markov models, or, in the case there are missing values, as hidden Markov models (HMM). A complication is that we aim to retain a traditional strength of duration analyses, namely, the ability to flexibly include covariates whose importance can be assessed by statistical tests, as in proportional hazard models (Box-Steffensmeier et al., 2007). A second inspiration of this work thus consists of using hidden Markov models with covariates, meaning that the transition rates of X and emission probabilities of Y can vary across individuals and time. This class of models, otherwise called Markov switching, has been regularly used in recent applied statistics works. Wang and Puterman (1999) used Markov Poisson regression models for discrete time series, and Langrock et al. (2017) proposed generalized additive models with B-splines for additional covariate flexibility. Recent applications include financial and economic forecasting (Meligkotsidou and Dellaportas, 2011), biometrics and environmental data (Wong, 2001), and shooting performance in sports (Sandri et al., 2020). These applications exhibit how Markov switching models produce good estimates yet remain easy to handle.
2.2.2 Formal presentation

### v3

# Some are observed, all leave traces: whole-population modeling of French elite civil servants’ career paths

@@ Abstract
Elite civil servants may come and go between the public and private sectors throughout their career, a process of particular interest for the public and social scientists. However, data to document such processes are rarely completely available: we need inference tools that can account for many missing values. We consider public-private paths of elite French civil servants and introduce binary Markov switching models with Bayesian data augmentation. Our procedure relies on two complementary data sources: (1) detailed observations of some individual trajectories obtained from LinkedIn; (2) less informative “traces” left by all individuals in the administrative record, which we model for missing data imputation. This model class maintains the properties of hidden Markov models and enables a tailored sampler to target the posterior, yet allows for varying parameters across individuals and time. By integrating the two sources, we can consider the whole population rather than just a sample, and avoid the biases that would stem from using only a single source. We demonstrate this allows to properly test substantive hypotheses on career paths across a variety of public organizations. We notably show that the probability for ENA graduates to exit the public sector has not increased since 1990, but that the probability they return has increased. We identify three clusters of organizations, with distinct patterns of public-private behaviors.
Keywords: Bayesian Markov Switching Model, Career Modeling, Public-Private Paths, Civil Servants, Missing Values, Data Augmentation
@@ 1 Introduction
Many important questions in the social sciences require understanding career trajectories of certain populations. This is an important aspect when describing political (jackle2018temporal), economic (dudouet_les_2010), academic (angermuller2023careers) and administrative (rouban_norme_2014) elites, but also when studying specific professions like musicians (abbott1990measuring), journalists (machut2023emploi), or even athletes (li2018journey). Studying careers is seen as interesting in its own right, but also to describe broader social mechanisms such as lobbying and revolving doors (blanes2012revolving) or the spatial structuring of finance (buhlmann2024career).
@@
However, a key problem in career modeling is data missingness. It is very rare that all data relevant to the career of a certain population are available in a public record or can even be collected because of resource constraints, which is problematic for two reasons. First, for a given defined population, inference can be made more difficult and less accurate if a majority of individuals cannot be observed. This makes the testing of interesting interaction effects impossible in practice. Second, researchers often choose to only study populations that can be entirely observed. This forces them to study the groups that are the most visible, such as top political elites. Both these issues thus constrain the breadth of questions that can be addressed using career modeling.
@@
Our focus is the study of public-private paths of elite civil servants in France, which have already been the subject of many quantitative publications (bouzidi_pantouflage_2010; rouban_norme_2014; kolopp2021pantoufler). Public-private paths broadly understood for policy-relevant positions are sometimes described using the term “revolving doors”, because these back-and-forth switches resemble a physical revolving door (seabrooke2021revolving; france_sphere_2017), but other scholars define revolving doors to be a much more precise and limited in scope phenomenon. Unassuming about this definitional debate, we denote this problem as the study of public-private paths of French elite civil servants, and acknowledge that our article falls in the first category of works.
Previous quantitative works of French civil servants have focused on data for a small number of individuals, typically collected by hand. Instead, our aim is to construct and apply a methodology to analyse the whole population, thanks to two complementary sources of data. First, it is relatively easy to describe the career paths of a portion of elite civil servants, by collecting data from digital networking websites such as LinkedIn on which they are registered. However, a majority of civil servants do not have a LinkedIn profile, and further, it is not always the case that a profile is fully completed from the starting point of one’s career. We thus complement the analysis with a second data source which covers the entire population over their entire public service career: to various degrees, civil servants leave “traces” of their activity, when they sign legal documents, or obtain important promotions or bonuses which are recorded in administrative records published in the Journal Officiel de la République Française (JORF). We describe these two data sources in detail below. Neither source taken alone can lead to robust inference on career paths: the LinkedIn data alone is insufficient as it concerns only a non-random subpopulation (in particular, individuals who have never left the public sector are less likely to be present on LinkedIn); the JORF data alone is insufficient as it contains too little information on its own.
We showcase that by combining these two sources and modeling them jointly, it is possible to reconstruct trajectories of large civil servant populations and draw interpretations that contribute to the social scientific literature: the union of the two data sources is greater than the sum of its parts. In addition, our population is more diverse and an order of magnitude larger compared to usual prosopography studies of the domain. Our methodology allows for analyses that are also targeted at less visible, larger professional groups. We describe our population of interest and data sources in Section 2.
In Section 3, we build hidden Markov models with covariates and autoregressive components and describe our inference algorithm. Section 4 showcases, using plausible synthetic data examples, that these components are required for such a modeling task while remaining simple enough for substantive hypothesis testing. We then apply these methods to data about French civil servants from 1990 to 2022 in Section 5, demonstrating the usefulness of this strategy and expanding the social scientific literature of administrative elites in meaningful ways. We consider in particular differences in behavior between organizations and through time.
Our methodology could plausibly be adapted to other the career analysis of other professional groups with dual data sources: one with high information about a minority of the population (possibly a hand curated data base), the other with low information traces on the totality of the population.
2 Problem setting
2.1 Population of interest
We consider elite French civil servants over the time frame 1990 to 2022. Top administrators are divided between groups and services that are sometimes tied to certain special statutes. While recent regulatory changes have attempted to increase mobility between them and foster a common organizational culture, they remain useful categories to base our approach upon.
Our study focuses on the following groups: three inspection groups whose mission is to audit public organizations organized thematically around financial affairs (igf, Inspection Générale des Finances), social affairs (igas, Inspection Générale des Affaires Sociales), and administration (iga, Inspection Générale de l’Administration); the Cour des Comptes (ccomptes), the supreme financial jurisdiction with a similar financial auditing function; the Conseil d’État (ce), whose members serve both as judges in the Supreme Administrative Court and as legal advisors to the government; the préfets (cprefet), state administration representatives in regions and departments; ambassadors (cdiplo, Corps Diplomatique); administrators of the national statistical and economics office (insee) who can serve key roles in the definition of economic policies; and finally, directors and assistant directors of two powerful services of the Ministry of the Economy, treasury (dgtresor), and public finances (dgfip). Additionally, we include in our study all graduates of the National School of Administration (École Nationale d’Administration, hereafter ENA). Our study includes all civil servants who joined or were promoted within one of these groups over the time frame 1990-2022.
These populations can be defined exhaustively, since all members of these groups are systematically described in the Journal Officiel de la République Française (JORF). This selection is voluntarily diverse and encompassing, as one of our goals is to describe how public-private mobility is a heterogeneous phenomenon. We obtained the complete list of population members from JORF, then used the LinkedIn API to find the profiles of these individuals. For each person in our population, we searched on LinkedIn for up to three profiles that share the same name, and then only kept the ones that list an experience in our starting set of organizations. For example, if an individual is included because they are part of the Statistical office, we only keep profiles with such an affiliation, using a rule-based approach. Overall, we designed our scheme to minimize false matchings between sources rather than maximizing the number of profiles that would be matched; manual inspection shows that it is very robust to homonymy problems in the matching between JORF and LinkedIn. The profiles output by the API queries were filtered using rules described in Appendix A, then all checked by hand.
Although the initial population is exhaustively defined, many individuals are not present on LinkedIn. The proportion of individuals with a LinkedIn match varies between groups, from 25% (cprefet) to 55% (ENA graduates); see Table 7 in Appendix B for all group proportions. Presence on LinkedIn is not independent of other quantities of interest. For example, Figure 1 shows that more information is available on digital profiles for the more recent period: the year of initial trace (a proxy for the year individuals joined the population of interest) is uniformly distributed across the whole time period, but the year of initial profile information (year from which each individual lists information on LinkedIn) is strongly biased towards more recent years, both because younger civil servants are more likely to have a LinkedIn profile and because older civil servants with a LinkedIn profile may only list their more recent positions. This is not the only bias in the LinkedIn data: for instance, we expect individuals who move to the private sector to be more likely to have a LinkedIn profile, compared to individuals who stay in the public sector for their entire career. Taken on their own, the LinkedIn data are thus not sufficient to examine the career paths of civil servants.
Figure 1: Comparison of the initial year where some information is available on an individual depending on whether it is the first administrative trace (left) or the first information on the digital profile (right).
We wish to understand the public-private paths at the level of the entire population, and to avoid biases that would stem from considering only individuals who are present on LinkedIn. The aim of this paper is to achieve this by augmenting the LinkedIn data with high coverage but low information data from JORF.
2.2 Previous works on career paths of French civil servants
Paths of civil servants in the private sector in France have been studied by multiple authors (bouzidi_pantouflage_2010; kolopp2021pantoufler; rouban_norme_2014) and have been a source of interest in the general public. We will explore two aspects assessed as interesting in the literature but not yet described, or not described with such precision: how the structure of public-private paths has evolved since 1990 (bouzidi_pantouflage_2010); and how patterns of these paths vary across many organizations (charle_pantouflage_1987) or individual characteristics such as gender (kolopp2021pantoufler). Previous studies have examined the trajectories of high civil servants in France by defining populations based on organizational belonging, but they have focused on detailed data sources only: well described career paths for a small number of people. As such, they have not explored thoroughly these questions, which require large amounts of data, and very diverse data.
rouban_inspection_2010 studied the finance general inspection (IGF) on the 1958-2008 period. IGF appears to have an important place in the revolving doors phenomenon, as they identified that of the 578 inspectors present in the organization between 1958-2009, a majority (62%) had wandered at least once in the private sector. They identify a discontinuity around 1990, with more individuals having attended a business school before joining ENA and then IGF. They also observed after 2000 a growing proportion of IGF members who have had an experience in the private sector before coming back to the public sector. However, this study is subject to bias because it does not account for the fact that different individuals have been observed over different time scales.
kolopp2021pantoufler conducted a study on the general direction of the treasury (dgtresor) in the 1965-2010 period. They exhibited a gendered behavior: men are more likely than women to leave for the most prestigious private industries (especially banking); on the other hand, they do not seem to show gender differences in the overall probability of leaving the public sector. They describe that differentiated practices generate unequal outcomes at different points in careers, for example with the stress on international mobility for career advancement and differences in professional networks. If we have reasons to assume these factors also play a significant role for other organizations, their strength may vary, and other factors might come into play. Even though we do not differentiate within private sector activities, we will compare behavior between men and women across the different organizations of our sample.
2.2.1 Analyses of ENA students
One subpopulation has been the subject of particular attention in the literature and in the public eye: former students of the National School of Administration ENA. ENA was founded in 1945 to select and train high civil servants for the French state. Graduates of this school (around 80 to 100 a year) play a central role in the political and administrative life of the country, and they constitute a natural population when studying elite civil servants.
rouban_norme_2014 conducted a survey over 6 student year-groups ($N=620$) who graduated in 1969-1970, 1989-90, and 1999-2000. He reports that, 10 years after graduating, 8.3% of students from the first two cohorts were working in a private company, compared to 14.5% for the two next cohorts, and 9.5% for the last two, which indicates year-to-year variation but not a clear trend towards increasing probability of leaving for the private sector. The results of rouban_norme_2014 seem to show that the average time taken for an ENA graduate before leaving the public sector has drastically decreased over time, which would imply a stark increase in the exit probability. We must however note that their methodology does not allow for comparing different cohorts. Indeed, many statistics considered in this study, like average time to leave to the private sector, are subject to right-censoring; this was not accounted for or acknowledged. This is problematic as the observation frames are significantly different across cohorts: it was, at the time of publication, 44 years for the 1970 cohort, but only 14 years for the 2000 cohort. Our results on similar questions in Section 5.2 leads to substantially different conclusions: we assume the difference with our results is due to our methodology properly taking into account the different lengths of observation frames. We conclude that the variation reported by rouban_norme_2014 is artefactual. The other quantitative conclusions of rouban_norme_2014 are not subject to this bias, and are coherent with our results.
bouzidi_pantouflage_2010 focus on a subgroup of ENA graduates, of the “civil administrator” status working in the Finance Ministry, between 1960 and 2002 ($N=438$), using hand-collected data. They show that, in their sample, about 40% of individuals went at least once into the private sector in the 20 years following graduation. The advantage of their approach is that they used a formal duration analysis model, which immunizes against right-censoring biases. They however do not attempt to estimate whether there was an acceleration or not on their period of study, and stress that doing such inference is difficult for later cohorts because of right-censoring. They also estimate time spent in the private sector using simulations using two Weilbull distributions, one for the public sector and one for the private sector. Our work can be seen as a discretization of such a procedure that also enables missing data imputation and inclusion of covariates.
2.3 Detailed and trace data sources
Unlike these previous analyses, we consider all individuals who joined or were promoted within the top French civil service at some point between 1990 and 2022. Some of these individuals stayed in the public sector for their whole career; others moved between the public and private sectors, possibly with multiple switches throughout their career. We wish to model, reconstruct, and understand these movements between public and private sectors.
@@ 2.3 Detailed and trace data sources
The core of our approach stems from the tension between data quality and quantity. In the majority of cases, it is difficult to obtain data that exhibit both high precision to document career trajectories, and high coverage. For French civil servants, we rely on two data sources with opposed strengths and weaknesses:
  • Detailed data source: LinkedIn. As mentioned above, we scraped the professional networking website LinkedIn for profiles of French civil servants; this gives us the exact path over some or all of these individuals’ careers. LinkedIn is a social media platform oriented towards professional networking, where users describe their past work experience on their profile. Many managers, both in the public and private sectors, have a profile, but not the entirety of the population. In addition, as these data are auto-biographical, individuals may choose to not include all relevant information: there are gaps in profiles, which means that there might be missing data even for individuals with a LinkedIn profile. This means that even when an individual is in this detailed data source, it can still be useful to complement the analysis with a trace data source. This data source exhibits low coverage but is very informative on the individuals it documents.
  • Trace data source: JORF. All civil servants leave traces in the Official Journal of the French Republic (JORF), an administrative record with a significant legal value that includes new pieces of regulation, executive decisions and nominations. This source documents nominations for many top-level positions in public organizations. It also includes, in certain cases, information about promotions, other career changes (such as retirement) and decisions taken by authorities that might involve named civil servants. Individuals thus leave traces both when they are nominated or promoted, and when they sign important administrative documents. Importantly, key positions are systematically described in JORF, and this source can thus be used to define an exhaustive population. However, not all public positions are described, and there is typically no duration attached to these positions, which explains why it is not sufficient taken on its own. JORF was digitized [^1] Data taken from Journal Officiel were first assembled and cleaned up by Nathann Cohen and put up on an API on the website steinertriples. Throughout the article, we rely on the amazing meta-data provided by this API, in addition to our own data cleaning procedures. starting in 1990, and this defines our temporal frame (1990-2022).
@@
    A typical trace in JORF might be a record of the form "on date $t$, individual $i$ was nominated to public position pp" or "on date $t$, individual $i$ signed a public document", both of which would show that $i$ worked for the public sector at date $t$. We usually do not have any information about when $i$ left their position, but we know that individuals in a certain career position should regularly leave traces of their activity in these records, although the rate at which this happens varies across positions and time. This data source exhibits high coverage but is not very informative taken on its own, and it requires statistical elaboration for analysis.
The LinkedIn data presents issues of heterogeneity, and data are missing not at random. We therefore expect any inference performed from the LinkedIn data alone to be biased; this is verified in practice in Section 5. On the other hand, the trace data source is exhaustive: traces are left for legal purposes, and depend only on the position occupied by a civil servant. It is however not possible to perform inference from the trace data alone, as the model parameters are not identifiable. Our modelling will therefore use both sources, and rely on the trace emission process to debias the inference. Concretely, it may be true that people on LinkedIn are more prone to go to the private sector, but if during their time in the public sector they leave traces in the same fashion as people who are not on LinkedIn (and after the effect of all regressors of the trace emission model are excluded), then adding trace data will be very effective at debiasing. There could still be some confounding effect because individuals with digital profiles do not occupy the same positions in the public sector than individuals not on online, and that these positions would be systematically linked to fewer declarations in JORF. It is not immediately clear what could lead to such an imbalance, and since this confounding would come from an average across many positions, it is reasonable to expect that it is a most of a much smaller order of magnitude than the bias that would result by only including digital profiles in the study.
This tension between complementary data sources could be found in other application areas. The detailed data source would typically correspond either to data collected by hand for a small number of individuals, or to data taken automatically from a source such as the professional networking website LinkedIn. The trace data source would correspond to larger administrative or digital records: for journalists, traces could be newspaper articles they signed; for researchers, traces would be academic papers; for the economic elite, traces might be having their names included in board meeting minutes. We focus on civil servants and do not explore these other potential applications, but we keep the methodological description general to facilitate its reuse.
3 Model building with Markov switching models and inference
3.1 Formalization
@@
We now turn to the formalization of this setting. We aim to study the career trajectory of $N$ individuals $i\in\{1,...,N\}$ in discrete time $t\in\{1,...,T\}$. We denote the career position of individual $i$ in time $t$ as $X_{i,t}\in\mathcal{X}$, with $\mathcal{X}=\{0,1\}$, corresponding to the public and private sectors; it would be straightforward to extend to a known finite state space $\mathcal{X}=\{x_{1},...,x_{K}\}$ of size $K$. The ensemble of all trajectories at all time points is denoted $\textbf{X}=(X_{i,t})_{i,t}$. Observe that time here is discrete: our techniques may be generalized with continuous time, but we do not find a sufficient interest given the prohibitive cost that would incur. We take $t$ to correspond to a six-month period. We assume that X is partially observed, thanks to the detailed data source: for some (but not all) values of $i$ and $t$, we observe the value of $X_{i,t}$.
@@
We now need to connect it to information found in the trace data source. The trace data source is described using a second set of time series, $\textbf{Y}=(Y_{i,t})_{i,t}$, which take value $Y_{i,t}\in\mathcal{Y}$ for each $i$ at time $t$. We take $Y_{i,t}$ to be a binary variable: $Y_{i,t}=1$ if an administrative trace can be found for individual $i$ at time $t$ and $Y_{i,t}=0$ otherwise. It would be straightforward to extend this to other discrete or continuous, one-dimensional or multi-dimensional spaces. Unlike X, we assume that Y is fully observed.
@@
Lastly, this setting is described with covariates $\textbf{W}=(W_{i,t})_{i,t}$ which are of interest to the researchers or which are deemed important to control for, and which may vary across individuals and time. A typical example will be to include the sex of individuals, their age, or their initial diplomas, to see if career behavior differs across these characteristics. Importantly, these covariates will be used both to model the behavior of career paths X, and the relationship between paths X and traces Y, and so may include both substantive and nuisance characteristics. We will show it is possible and important to include transformations of lagged values of $Y_{i,t}$ into $W_{i,t}$ when describing the relationship between X and Y.
@@
It is natural to study discrete trajectories as Markov models, or, in the case there are missing values, as hidden Markov models (HMM). A complication is that we aim to retain a traditional strength of duration analyses, namely, the ability to flexibly include covariates whose importance can be assessed by statistical tests, as in proportional hazard models (box-steffensmeier_event_2007). We thus consider hidden Markov models with covariates, meaning that the transition probabilities of X and emission distributions of Y can vary across individuals and time. Such models are called Markov switching models. They produce good estimates yet remain easy to handle, and have been used across a wide range of applications such as clinical trials (wang_markov_1999), finance (langrock_markov-switching_2017), biometrics and environmental data (wong_logistic_2001), and shooting performance in sports (sandri_markov_2020).
3.2 Formal presentation